OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x − 9

OpenStudy (eyust707):

Here we need to bring all the terms with x's in them to one side and all the terms with numbers in them to the other side.

OpenStudy (ghazi):

5x-4x=-9-2 x=-11